OpenStudy (anonymous):

Write in point slope form, the equation of a line with a slope of -3 that goes through point (-4,5)

OpenStudy (anonymous):

y-y1=m(x-x1) y1=5, m=-3, x1=-4 y-5= -3(x+4) >>>point slope y-5=3x-12>>add 5 both sides y=3x-7 >>>slope intercept